OpenStudy (anonymous):

Medal to most helpful. Short 2 part question. 1. Which of the following statements have the same result? 2. Explain each step in solving each one.

OpenStudy (anonymous):

I. f(2) when f(x) = 3x + 2 II. \[f^{-1}(3) when f(x)=\frac{ 2x-7 }{ 3 }\] III. 2y + 14 = 4y − 2

OpenStudy (anonymous):

Work for I: f(x) = 3x + 2 f(2) = 3(2) + 2 = 6 + 2 = 8 Work for III: 2y + 14 = 4y − 2 -2y - 2y 14 = 2y - 2 + 2 + 2 16 = 2y y = 8 So, I and III both equal 8. My problem is that I don't know how to figure out II. Help!?

OpenStudy (anonymous):

Will you help me out, please? @Hero @Abhisar @nincompoop @jim_thompson5910 @mathmale @SolomonZelman @undeadknight26

OpenStudy (anonymous):

maybe this will help http://openstudy.com/study#/updates/54667d7ae4b05b0e1f1d40b3

OpenStudy (anonymous):

Thanks.